# Selecting informative conformal prediction sets with false coverage rate control

### v1

@@ Abstract
In supervised learning, including regression and classification, conformal methods provide prediction sets for the outcome/label with finite sample coverage for any machine learning predictors. We consider here the case where such prediction sets come after a selection process. The selection process requires that the selected prediction sets be ‘informative’ in a well defined sense. We consider both the classification and regression settings where the analyst may consider as informative only the sample with prediction label sets or prediction intervals small enough, excluding null values, or obeying other appropriate ‘monotone’ constraints. While this covers many settings of possible interest in various applications, we develop a unified framework for building such informative conformal prediction sets while controlling the false coverage rate (FCR) on the selected sample. While conformal prediction sets after selection have been the focus of much recent literature in the field, the new introduced procedures, called InfoSP and InfoSCOP, are to our knowledge the first ones providing FCR control for informative prediction sets. We show the usefulness of our resulting procedures on real and simulated data.
Keywords: conformal prediction, selective inference, false coverage rate, label shift, classification, regression, prediction interval.
@@ 1 Introduction
1.1 Background
In modern data analysis, machine learning algorithms are often used to make predictions and a main challenge is to measure the uncertainty of such methods. Conformal inference offers an elegant solution to this problem, by providing prediction sets that provably cover the true value with high probability, for any sample size, any predictive algorithm, and any distribution of the data Vovk et al., (2005). We consider the following classic split/inductive conformal prediction setting Papadopoulos et al., (2002); Vovk et al., (2005); Lei et al., (2014). Let $(X,Y)\in\mathcal{X}\times\mathcal{Y}$ be a random vector with unknown distribution $P_{XY}$. Typically, $\mathcal{X}$ is a subset of $\mathbb{R}^{d}$ (real valued covariates), and $\mathcal{Y}=[K]$ (classification of among $K\geq 2$ classes) or $\mathcal{Y}=\mathbb{R}$ (regression, with a real valued outcome) [^1] Our theory can be applied for more general observation spaces (e.g., $\mathcal{Y}=\mathbb{R}^{d^{\prime}}$) but we consider the most common settings $\mathcal{Y}\in\{[K],\mathbb{R}\}$ for simplicity.. In this setting, there are two independent samples of points $(X_{i},Y_{i})$: the calibration sample $\{(X_{j},Y_{j}),j\in[n]\}$; and the test sample $\{(X_{n+i},Y_{n+i}),i\in[m]\}$.
While all measurements are observed in the calibration sample, only the $X_{i}$’s are observed in the test sample and the aim is to provide prediction sets for the unobserved $(Y_{n+i},i\in[m])$. We denote the prediction set for $Y_{n+i}$ by $\mathcal{C}_{n+i}$. It is a subset of $[K]$ for classification and a subset (often an interval) of $\mathbb{R}$ for regression. The classic conformal prediction set for $Y_{n+i}$, denoted by $\mathcal{C}^{\alpha}_{n+i}$, is function of $\{(X_{j},Y_{j}),j\in[n]\},X_{n+i}$, and $\alpha$ (Sadinle et al.,, 2019; Lei et al.,, 2018) and has the following coverage guarantee, assuming that $(X_{n+i},Y_{n+i})$ and $(X_{j},Y_{j}),j\in[n]$, are exchangeable pairs of observations from any distribution $P_{XY}$:
@@ 1.1 Background
In practice, the size of the test sample $m$ is often large, encompassing hundreds or thousands of unlabeled examples. Inferring on all of them is unnecessary or inefficient in many applications (Jin and Candes,, 2023; Bao et al.,, 2024). For example, in classification, if each image belongs to one of $[K]$ categories, the analyst is not interested in the examples where $\mathcal{C}^{\alpha}_{n+i}=[K]$. Thus, it is natural to assume that a subset of individuals will be selected. However, reporting their prediction sets constructed to have at least $1-\alpha$ confidence is problematic, since conditional on being selected, the coverage may be much smaller (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013).
1.2 Informative prediction sets
@@
    Length-restricted intervals: =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>|I|\leq 2\lambda_{0}\}$ for some size $\lambda_{0}>0$. It is useful if the user only wants to report prediction intervals that are accurate enough.
@@
    Non-trivial classification: =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq K-1\}$. It is appropriate when the user wants a label set that is minimally informative. More generally, at most $k_{0}$-sized classification can be considered with =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq k_{0}\}$.
A common target error guarantee is that the inference on at most $\alpha$ examples among the selected is expected to be false. This is a classical error criterion in the selective inference literature, (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013). It has been used, e.g., for novelty detection (Bates et al.,, 2021; Marandon et al.,, 2024), for classification (Rava et al.,, 2021; Zhao and Su,, 2023; Jin and Candes,, 2023), for regression (Bao et al.,, 2024), and for unsupervised clustering (Mary-Huard et al.,, 2022; Marandon et al.,, 2022). For selecting prediction sets, the target error guarantee is thus that at most $\alpha$ examples among the selected are expected to have prediction sets that do not cover their true outcome value. The false coverage proportion (FCP) for the procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ is defined as the proportion of non-covered examples in the selected set $\mathcal{S}$:
@@
In this work, we consider two popular models for generating the samples [^2] In both models, the independence assumption can be relaxed. In the iid model, it is enough that all $n+m$ samples are exchangeable. In the class conditional model, it is enough that the subset of $[n+m]$ of samples from the same class is exchangeable, for all classes in $[K]$. See § A for more details.:
  • the iid model (both for regression and classification): the variables $(X_{i},Y_{i})\sim P_{XY}$, $i\in[n+m]$, are all independent and identically distributed (iid). This is the standard assumption classically used for conformal prediction (Vovk et al.,, 2005). The parameter of the model is in that case $P_{XY}$;
@@
  • the class-conditional model (only the classification setting): all the labels $(Y_{i},i\in[n+m])$ are deterministic and the covariates $(X_{i},i\in[n+m])$ are mutually independent with $X_{j}\sim P_{X|Y=Y_{j}}$. It relaxes the exchangeable model assumptions of iid model, by only requiring that the distribution within each class is the same for the test and calibration sample (Sadinle et al.,, 2019; Ding et al.,, 2023). The target FCR is then conditional on the labels. The parameters of the model are then given by $(P_{X|Y=k})_{k\in[K]}$ and $(Y_{i},i\in[n+m])$.
1.3 Contributions
We briefly summarize here the contributions of our work.
@@ 1.3 Contributions
  • We first introduce a new method, called InfoSP (Informative selective prediction set), that selects only informative prediction sets with a level $\alpha$ FCR guarantee on the selected (§ 3.1). Formally, this means that we achieve both $\mathrm{FCR}(\mathcal{R})\leq\alpha$ and $\forall i\in\mathcal{S}$, $\mathcal{C}_{n+i}\in\mathcal{I}$, for $\mathcal{I}$ being the collection of informative subsets. The FCR control of InfoSP is established both in the iid model and class-conditional model (see Theorem 3.1).
@@
  • We introduce a second procedure, called InfoSCOP (Informative selective conditional prediction set), that has the same theoretical properties as InfoSP in the iid model (§ 3.2) and that relies on an initial selection step that is aimed at eliminating (at least some of) the examples for which informative prediction sets cannot be constructed. Further selection then takes place in order to ensure that all reported prediction sets are informative. While the pre-processing step is inspired by Bao et al., (2024), the theoretical framework in Bao et al., (2024) precludes this type of selection (see § 1.4 for more details).
@@
  • Our main theoretical FCR control guarantees come from a single general theorem in § A.1 that generalizes the theoretical results in Benjamini and Yekutieli, (2005); Benjamini and Bogomolov, (2013), to include more general classes of $p$-values and selection rules. This goes in the direction of unifying theory for selective conformal prediction sets.
  • We optimize the analysis pipeline for common informative selection rules in § 4 (regression) and § 5-§ 7 (classification), while providing additional theoretical results and appropriate numerical experiments.
@@
To provide an intuition for our approach, InfoSP is illustrated in Figure 1 for the regression case and in Figure 2 for the classification case (for $K=3$ classes). Left panels display a naive method reporting the marginal classical conformal prediction sets $\mathcal{C}^{\alpha}_{n+i}$ in (1) for all those that are informative, that is, such that $\mathcal{C}^{\alpha}_{n+i}\in\mathcal{I}$, without further correction. As expected, this method inflates the FCR values above $\alpha$ and thus does not provide the desired FCR control. The new procedure InfoSP is displayed in the right panels: we can see that only informative prediction intervals are selected while FCR is maintained close to $\alpha$, thanks to a suitable correction in the prediction sets.
Classical Conformal InfoSP
@@
Refer to caption Refer to caption
Figure 1: Informative prediction intervals in regression (iid setting). First row: when excluding $[a,b]$ (Example 1.1 item 1) in an homoscedastic Gaussian regression model with perfect variance prediction. Second row: length-restricted (Example 1.1 item 2) in an heteroscedastic Gaussian regression model with perfect mean prediction. $\alpha=10\%$. These panels are extracted from Figures 3 and 4, see therein for more details.
Classical Conformal InfoSP
(FCR = 0.160) (FCR = 0.099)
Refer to caption Refer to caption
@@
Figure 2: Informative prediction sets in classification for CIFAR-10 dataset, restricted to the $K=3$ classes “bird”, “cat”, and “dog” classes (iid setting). Informative prediction subsets are those of size smaller than $K-1=2$ (i.e., non-trivial, Example 1.2 item 1). Selection by InfoSP are framed in red (right panel). $\alpha=10\%$. See § 7 for more details.
1.4 Connections to existing works
We describe here inspiring works that are the most connected to ours. All these works assume the iid model.
Zhao and Su, (2023) suggested procedures for average error control in multi-class classification, so $\mathcal{C}_{n+i}$ are singletons. Since their procedure only reports a single class for each selected example, ambiguous examples will not be selected. However, in most classification tasks, there are examples whose true class is difficult to determine, yet it is possible to narrow down the possible set of classes (Sadinle et al.,, 2019). We suggest procedures that produce $\mathcal{C}_{n+i}$ that are not necessarily singletons for $K>2$. For $K=2$, their procedure coincides with an instance in our framework, see details in § 5.4. However, for $K>2$, while our suggestion as well as their suggestion provides level $\alpha$ FCR control, we select more examples, and although the prediction sets may be at a coarser resolution than singletons, they are still informative since they narrow down the possible set of classes.
Jin and Candes, (2023) addressed the problem of discovering outcomes with values above a threshold. So $\mathcal{C}_{n+i}$ is of the form $(c_{i},\infty)$ for predefined $(c_{i})_{i\in[m]}$. They cast the problem as that of testing the family of null hypotheses $\{Y_{n+i}\leq c_{i},i\in[m]\}$. In § E.2, it is demonstrated that by defining $Y_{n+i}\leq c_{i}$ as uninformative, we can complement the discoveries of Jin and Candes, (2023) with one-sided prediction intervals, while providing the same false discovery rate guarantee on the selected. Moreover, we show how to obtain two-sided prediction intervals for the informative examples in § 4.1.
Bao et al., (2024) considered the regression framework. Their first result (Proposition 1 therein) is to prove that for selection rules that do not depend on the calibration sample, classic conformal prediction intervals at level $\alpha|\mathcal{S}|/m$ for the $\mathcal{S}$ selected examples (i.e., the correction factor $|\mathcal{S}|/m$ for selection suggested in Benjamini and Yekutieli,, 2005), provide level $\alpha$ FCR control. For our purpose of informative selection, this result is not useful because informative selection involves all conformal $p$-values and therefore involves the calibration sample in a specific way. We need a different set of assumptions that are detailed in our novel Theorem A.1.
Bao et al., (2024) further argue that the resulting prediction intervals, $\big{(}\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}\big{)}_{i\in\mathcal{S}}$ are too wide. They suggest a novel approach that performs selection on both the calibration set and test set, and then constructs $\alpha$ level conformal prediction intervals for the selected test candidates using the conditional empirical distribution obtained by the post-calibration set. For exchangeable selection rules, they show that the FCR is controlled at level $\alpha$. Their selection process cannot guarantee that all the constructed prediction intervals are of interest to the analyst. For example, for predicting the affinity of drug-target pairs, the analyst may not be interested in pairs with affinity below, say, $y_{0}$ (the case in item 1 of Example 1.1). Using the novel approach of Bao et al., (2024), prediction intervals will be constructed following the selection of calibration and test examples for which the predicted affinity from the machine learning algorithm is above a selection threshold. They require that the selection procedure be a thresholding procedure of the scores $S_{Y_{j}}(X_{j})$ with a threshold $\tau$ that is either independent of the calibration sample (their Proposition 1), or exchangeable with respect to both calibration and test samples (their Theorem 1). Some of these prediction intervals may include $y_{0}$, and thus be useless for the analyst (an illustration is given in § E.3 in SM, see Figure 12 therein). However, it is not possible to additionally select only the examples with prediction intervals above $y_{0}$, since after performing this additional selection, the prediction intervals of Bao et al., (2024) on the selected no longer have an $\alpha$ level FCR guarantee. Our procedures for regression thus complement the work of Bao et al., (2024) when the focus is that all the prediction intervals eventually constructed are informative.
In a very recent work [^3] This work appeared when we were in the final stage of writing, our work has been done independently., Jin and Ren, (2024) generalize the work of Bao et al., (2024), by considering more general selection rules for finite-sample exact coverage conditional on the unit being selected. Their conditional guarantee is achieved by a careful swapping argument which identifies for each selection rule, the appropriate subset of the calibration examples for each example from the test sample. Their conditional error guarantee implies FCR control under some conditions. They provide various examples where selection takes place and prediction sets on the selected are of interest. Even though their selection may be based on some notion of informativeness, like in Bao et al., (2024) the final prediction sets that are constructed can violate this notion. For example, after selecting by a multiple testing procedure on the family of null hypotheses $\{Y_{n+i}\leq c_{i},i\in[m]\}$ in the setting of Jin and Candes, (2023), their prediction sets may include the $c_{i}$’s for some of the discoveries. We suggest procedures where selection and construction of prediction sets are inseparable, since we require that each selected prediction set be informative (along with the requirement that FCR $\leq\alpha$).
Notation
Expectations and probabilities are denoted for the iid model with $\mathbb{E}_{(X,Y)\sim P_{XY}}(\cdot)$ and $\mathbb{P}_{(X,Y)\sim P_{XY}}(\cdot)$, and for the class-conditional model with $\mathbb{E}_{X\sim P_{X|Y}}(\cdot)$ and $\mathbb{P}_{X\sim P_{X|Y}}(\cdot)$, respectively. If the data generation process is clear from the context, or if the expression is relevant for both models, then the subscript is omitted. In addition, $A\subset B$ means that the set $A$ is included in the set $B$ with a possible equality. Finally, for any subset $C\subset\mathbb{R}$, $|C|$ denotes the cardinality of $C$ if the set $C$ is finite, and the Lebesgue measure of $C$ if $C$ is an interval.

### v3

@@ Abstract
In supervised learning, including regression and classification, conformal methods provide prediction sets for the outcome/label with finite sample coverage for any machine learning predictor. We consider here the case where such prediction sets come after a selection process. The selection process requires that the selected prediction sets be ‘informative’ in a well defined sense. We consider both the classification and regression settings where the analyst may consider as informative only the sample with prediction sets small enough, excluding null values, or obeying other appropriate ‘monotone’ constraints. We develop a unified framework for building such informative conformal prediction sets while controlling the false coverage rate (FCR) on the selected sample. While conformal prediction sets after selection have been the focus of much recent literature in the field, the new introduced procedures, called InfoSP and InfoSCOP, are to our knowledge the first ones providing FCR control for informative prediction sets. We show the usefulness of our resulting procedures on real and simulated data.
Classification,
false discovery rate,
label shift,
prediction interval,
regression,
selective inference,
keywords:
@@ 1 Introduction
In modern data analysis, machine learning algorithms are often used to make predictions and a main challenge is to measure the uncertainty of such methods. Conformal inference offers an elegant solution to this problem, by providing prediction sets that provably cover the true value with high probability, for any sample size, any predictive algorithm, and any distribution of the data Vovk et al., (2005). We consider the following classic split/inductive conformal prediction setting Papadopoulos et al., (2002); Vovk et al., (2005); Lei et al., (2014). Let $(X,Y)\in\mathcal{X}\times\mathcal{Y}$ be a random vector with unknown distribution $P_{XY}$. Typically, $\mathcal{X}$ is a subset of $\mathbb{R}^{d}$ (real valued covariates), and $\mathcal{Y}=[K]$ (classification among $K\geq 2$ classes) or $\mathcal{Y}=\mathbb{R}$ (regression, with a real valued outcome) [^1] Our theory can be applied for more general observation spaces (e.g., $\mathcal{Y}=\mathbb{R}^{d^{\prime}}$) but we consider the most common settings $\mathcal{Y}\in\{[K],\mathbb{R}\}$ for simplicity.. In this setting, there are two independent samples of points $(X_{i},Y_{i})$: the calibration sample $\{(X_{j},Y_{j}),j\in[n]\}$; and the test sample $\{(X_{n+i},Y_{n+i}),i\in[m]\}$. We also have a prediction machine for $Y_{j}$ given $X_{j}$, built from training data that is independent of the calibration and test data.
While all measurements are observed in the calibration sample, only the $X_{i}$’s are observed in the test sample and the aim is to provide prediction sets for the unobserved $(Y_{n+i},i\in[m])$. We denote the prediction set for $Y_{n+i}$ by $\mathcal{C}_{n+i}$. It is a subset of $[K]$ for classification and a subset (often an interval) of $\mathbb{R}$ for regression. The classic conformal prediction set for $Y_{n+i}$, denoted by $\mathcal{C}^{\alpha}_{n+i}$, is a function of $\{(X_{j},Y_{j}),j\in[n]\},X_{n+i}$, and $\alpha$ (Sadinle et al.,, 2019; Lei et al.,, 2018) and has the following coverage guarantee, assuming that $(X_{n+i},Y_{n+i})$ and $(X_{j},Y_{j}),j\in[n]$, are exchangeable pairs of observations from any distribution $P_{XY}$:
@@
In practice, the size of the test sample $m$ is often large, encompassing hundreds or thousands of unlabeled examples. Inferring on all of them is unnecessary or inefficient in many applications (Jin and Candes,, 2023; Bao et al.,, 2024). For example, in classification, if each image belongs to one of $[K]$ categories, the analyst is not interested in the examples where $\mathcal{C}^{\alpha}_{n+i}=[K]$. Thus, it is natural to assume that a subset of individuals will be selected. However, reporting their prediction sets constructed to have at least $1-\alpha$ confidence is problematic, since conditional on being selected, the coverage may be much smaller (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013), see also Figure 1 below.
@@
    Length-restricted intervals: =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>|I|\leq 2\lambda_{0}\}$ for some $\lambda_{0}>0$, which are useful for only reporting prediction intervals that are accurate enough.
@@
    Non-trivial classification: =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq K-1\}$. It is appropriate when the analyst wants a label set that is minimally informative. More generally, at most $k_{0}$-sized classification can be considered with =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq k_{0}\}$.
A common target error guarantee is that the inference on at most $\alpha$ examples among the selected is expected to be false. This is a classical error criterion in the selective inference literature, (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013; Weinstein and Ramdas,, 2020). It has been used, e.g., for novelty detection (Bates et al.,, 2023; Marandon et al.,, 2024), for classification (Rava et al.,, 2021; Zhao and Su,, 2023; Jin and Candes,, 2023), for regression (Bao et al.,, 2024), and for unsupervised clustering (Mary-Huard et al.,, 2022; Marandon et al.,, 2022). For selecting prediction sets, the target error guarantee is thus that at most $\alpha$ examples among the selected are expected to have prediction sets that do not cover their true outcome value. The false coverage proportion (FCP) for the procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ is defined as the proportion of non-covered examples in the selected set $\mathcal{S}$:
@@
In this work, we consider two popular models for generating the samples [^2] In both models, the independence assumption can be relaxed. In the iid model, it is enough that all $n+m$ samples are exchangeable. In the class-conditional model, it is enough that the subset of $[n+m]$ of samples from the same class is exchangeable, for all classes in $[K]$. See § C for more details.. First, the iid model (both for regression and classification): the variables $(X_{i},Y_{i})\sim P_{XY}$, $i\in[n+m]$, are all independent and identically distributed (iid). This is the standard assumption classically used for conformal prediction (Vovk et al.,, 2005). The parameter of the model is in that case $P_{XY}$. Second, the class-conditional model (only the classification setting): all the labels $(Y_{i},i\in[n+m])$ are deterministic and the covariates $(X_{i},i\in[n+m])$ are mutually independent with $X_{j}\sim P_{X|Y=Y_{j}}$. It relaxes the exchangeable model assumptions of iid model, by only requiring that the distribution within each class is the same for the test and calibration sample (Sadinle et al.,, 2019; Ding et al.,, 2023). The target FCR is then conditional on the labels. The parameters of the model are then given by $(P_{X|Y=k})_{k\in[K]}$ and $(Y_{i},i\in[n+m])$.
@@
We now briefly summarize the contributions of our work. We first introduce a new method, called InfoSP (Informative selective prediction sets), that selects only informative prediction sets with a level $\alpha$ FCR guarantee on the selected (§ 3.1). Formally, this means that we achieve both $\mathrm{FCR}(\mathcal{R})\leq\alpha$ and $\forall i\in\mathcal{S}$, $\mathcal{C}_{n+i}\in\mathcal{I}$, for $\mathcal{I}$ being the collection of informative subsets. The FCR control of InfoSP is established both in the iid model and class-conditional model (see Theorem 3.1).
@@
We introduce a second procedure, called InfoSCOP (Informative selective conditional prediction sets), that has the same theoretical properties as InfoSP in the iid model (§ 3.2) and that relies on an initial selection step that is aimed at eliminating (at least some of) the examples for which informative prediction sets cannot be constructed. Further selection then takes place in order to ensure that all reported prediction sets are informative. While the pre-processing step is inspired by Bao et al., (2024), their theoretical framework precludes this type of selection (see § A for more details).
@@
Third, our main theoretical FCR control guarantees come from a single general theorem in § C.1 that includes general classes of $p$-values and accommodates any concordant selection rule (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013). Importantly, our novel theory supports conformal $p$-values, and selecting only informative prediction sets, as specific examples.
We optimize the analysis pipeline for common informative selection rules in § 4 (regression) and § 5-§ 6 (classification), while providing additional theoretical results and appropriate numerical experiments. The interest of the method is also demonstrated in a specific application in the field of molecular biology in § I.3, for which the practitioner aims at predicting gene expression from promoter sequences. While the analyst is interested in selecting promoters that cause low-expression, the produced prediction two-sided interval allows to establish that the expression is above zero and to evaluate the plausible extent to which it exceeds zero. Finally, an application to directional FDR control is also provided in § B.
@@
To provide an intuition for our approach, InfoSP is illustrated in Figure 1 for the classification case (for $K=3$ classes). Left panels display a naive method reporting the marginal classical conformal prediction sets $\mathcal{C}^{\alpha}_{n+i}$ in (1) for all those that are informative, that is, such that $\mathcal{C}^{\alpha}_{n+i}\in\mathcal{I}$, without further correction. Since no error can occur when the prediction set is trivial, the selection always inflates the FCP values. The new procedure InfoSP is displayed in the right panels: prediction sets are made slightly larger to maintain a correct FCP value while the selection (red boxes) guarantees that only informative (i.e., non-trivial) prediction sets are selected. This example is only for one data generation: it is presented here only for illustrative propose and more accurate in-expectation FCR values are given in § 6. The regression case is illustrated in Figures 2 and 3 (§ 4), for which the second procedure InfoSCOP is also displayed. It is worth to note that in some situations, the latter may even results in prediction sets that are smaller (!) than those of the naive method.
1.1 Relation to previous work
There are interesting connections between our approach and previous work of the literature: selecting confidence interval that satisfies specific notions of informativeness Weinstein and Yekutieli, (2020); Weinstein and Ramdas, (2020); multi-class classification Zhao and Su, (2023); and very recent works on selective conformal inference Bao et al., (2024); Jin and Ren, (2024). Due to the limited space, the details are in§ A of the SM.
We also underline that our second procedure InfoSCOP relies on the approach of splitting the calibration sample, which has already been used in conformal literature in different contexts. The idea is to enable an additional data-driven tuning of the method by only paying the price of splitting the calibration sample. For instance, it has been provided in Marandon et al., (2024) for the task of building adaptive scores. In the present aim of controlling the FCR on a data-driven selection, we formulate a general statement in the SM, see Lemma H.5. It applies to any type of data-driven selection and error rate.
Classical conformal with naive selection InfoSP
(FCP = 3/20 = 0.15) (FCP = 1/10 = 0.10)
@@ 1.1 Relation to previous work
Figure 1: Informative prediction sets in classification for CIFAR-10 dataset, restricted to the $K=3$ classes “bird”, “cat”, and “dog” classes (iid setting). Informative prediction subsets are those of size at most $K-1=2$ (i.e., non-trivial, Example 1.2 item 1). Selection by InfoSP are framed in red (right panel). $\alpha=10\%$. See § 6 for more details.